the average of four numbers is 638.4 the first four numbers are 216 159.8 and 301.25 Find the fourth number.: the average of four numbers is 638.4 the first four numbers are 216 159.8 and 301.25 Find the fourth number.
Answer by edjones(2415)

216+159.8+301.25+x=638.4*4
677.05+x=2553.6
x=2553.6-677.05
x=1876.55
.
Ed
Question 163277: the average of four numbers is 638.4 the first four numbers are 216 159.8 and 301.25 Find the fourth number.: the average of four numbers is 638.4 the first four numbers are 216 159.8 and 301.25 Find the fourth number.
Answer by orca(336)

Let the fourth number be x.
Then the average of the four number can be expressed as:
(216 + 159.8 + 301.25 + x)/4
As the average is 638.4, we have
(216 + 159.8 + 301.25 + x)/4=638.4
Solving the equation for x, we obtain
(677.05 + x)/4=638.4
677.05 + x=2553.6 (multiply both sides by 4)
x=1876.55